OpenStudy (anonymous):

5 Algebra 2, multiple choice questions? Please help me solve for x....

OpenStudy (anonymous):

\[\frac{ 3 }{ 3x }+\frac{ 1 }{ x+4 }=\frac{ 10 }{ 7x }\]

OpenStudy (anonymous):

A. x=-3 B. x=3 C. x=-3 and x=3 D. No Solution

OpenStudy (anonymous):

\[\frac{ 5 }{ x^2-1 }-\frac{ 2 }{ x }=\frac{ 2 }{ x+1 }\] A. x=1/4 and 2 B. x=-1/4 and -2 C. x=1/4 and -2 D. x=-1/4 and 2

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 }-\frac{ 1 }{ x }=\frac{ 3 }{ 4 }\] A. x= 1 B. x=2 C. x=4 D. x=8

OpenStudy (anonymous):

\[\frac{ 2 }{ 2 }+\frac{ 4 }{ 2x }=\frac{ x+4 }{ 10 }\] A. x=-4 B/ x=5 C. x=-4 and x= 5 D. No Solution

OpenStudy (anonymous):

\[\frac{ 5 }{ x }+\frac{ 4 }{ x^-9 }=\frac{ 3 }{ x-3 }\] A. x=-4 and x=-9 B. x=4 and x=-9 C. x=-4 and x=9 D. x=4 and x=9

OpenStudy (amistre64):

multiple choice, if stuck, simply amounts to testing if the options fit the bill

OpenStudy (anonymous):

Plug in the values for x to See what matches?

OpenStudy (amistre64):

yep ... that pretty much simplifies things to addition and subtraction of fractions in this case

OpenStudy (anonymous):

Okay, thanks :) I'll give it a try

OpenStudy (amistre64):

\[\frac{ 5 }{ x }+\frac{ 4 }{ x^{-9} }=\frac{ 3 }{ x-3 }\]does x=4 ring true? \[\frac{ 5 }{ 4 }+\frac{ 4 }{ 4^{-9} }=\frac{ 3 }{ 4-3 }\] \[\frac{ 5 }{ 4 }+4^{10} =3\]im going with x not equal 4 in that case ... so we can rule out that option

OpenStudy (amistre64):

that leaves options A and C, since they differ by -9 or 9 try the x=9 option

OpenStudy (amistre64):

if it fits, your good, if not ... you know its the other one

OpenStudy (anonymous):

Process of elimination :)

OpenStudy (amistre64):

exactly ;) thats how i passed most of my community college tests

OpenStudy (anonymous):

Haha great advice :)

OpenStudy (amistre64):

check that x^(-9) i think youve got that typoed ...

OpenStudy (anonymous):

Yes. It should be x-9 not exponents

OpenStudy (anonymous):

#1=B

OpenStudy (anonymous):

#2=D

OpenStudy (amistre64):

#1 is b, correct

OpenStudy (anonymous):

Thanks :)

OpenStudy (amistre64):

#2 is d, correct

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

#3=D

OpenStudy (amistre64):

correct, that one was rather simple 3/4 = 6/8 7/8 - 1/8 = 6/8 ... so x=8

OpenStudy (amistre64):

i think the last one is still typoed; you might wanna rewrite it

OpenStudy (anonymous):

#4=A

OpenStudy (anonymous):

#5=C

OpenStudy (anonymous):

#4 is not A

OpenStudy (amistre64):

it looks like your on a roll :) what is the problem setup for #4?

OpenStudy (anonymous):

When I plugged in -4 for x in #4, it worked. But when I submitted my answer it was wrong. I should have worked all of them before submitting :( Oh, well :) now I know for next time. That was the only one I got wrong

OpenStudy (amistre64):

ahh, well now you also know the shortcut ;) good job overall

OpenStudy (anonymous):

Thanks! :D